What is the probability that 2 (or more) people in a random group of 25 people have a common birthday? (This is the famous Birthday Paradox Problem.)
Answer by sudhanshu_kmr(1152)


Here we firstly find the probability of event when all 25 people have different birthday...
probability of first people have its own birthday = 365/365 ( sure )
probability of second people have different b'day to first = 364/365

probability of third people have different b'day to first and second = 363/365
:
:
:
:
similarly for 25th person, probability of different b'day = 341/365

probability when all have different b'day P(A) = 365/365 * 364/365 *.....*341/365

= (1/365)^24 * ( 364 *.....*341)

now, probability that 2 or more people have common b'day = P(A')

= 1-P(A)